OpenStudy (anonymous):

rewrite the equation below in standard form. 2/3x=-5/2y+2 A. 4x+15y=12 B. 2x+15/2y=6 C. 2x+5y=4 D. 2x+5y=6

OpenStudy (jim766):

first lets get rid of the fractions \[\frac{ 2x }{ 3 } = \frac{ -5y }{ 2 }+2\] mult everything by common denominator 6 \[\frac{ 6(2x) }{ 3 } = \frac{6( -5y) }{ 2 }+6(2)\] now simplify \[\frac{ 12x }{ 3 } = \frac{-30y }{ 2 }+12\] simplify more \[4x = -15y + 12\] now need to get x and y on same side 4x = -15y + 12 +15y +15y 4x + 15y = 12 no fractions, x coefficient is positive, now common factor